### v1

# Bayesian Constraints on Pre-Equilibrium Jet Quenching and Predictions for Oxygen Collisions

Abstract
The contrast between the as-yet unmeasurable energy-loss effects in proton-nucleus collisions and the striking magnitude of the so-called high-momentum flow coefficients challenges our understanding of jet quenching mechanisms in large nucleus-nucleus collisions when applied to smaller systems. Intermediate-sized, light ion collisions will offer key insight into the system-size dependence of the interplay between jet energy loss and jet flow effects. To make quantitative predictions, we extend a semi-analytic jet quenching framework by coupling it to state-of-the-art event-by-event hydrodynamics and, for the first time, incorporate pre-equilibrium energy loss via the hydrodynamic attractor. A Bayesian analysis shows that an early-time onset of energy loss is compatible with RHIC and LHC measurements of jet suppression and jet elliptic flow in large systems, as well as hadron suppression, with the exception of hadron elliptic flow. Using these constraints, we predict both hadron and jet quenching observables in oxygen-oxygen collisions, finding sizable energy loss that exceeds the no-quenching baseline.
I Introduction
Understanding the emergence of collective behavior in small collision systems, such as $pp$ and $pA$ collisions, remains one of the central challenges for both the high-energy physics and heavy-ion communities since their discovery at the LHC more than a decade ago [1, 2, 3, 4] (for recent reviews, see [5, 6, 7]). While the dynamical origin of this collectivity is still under investigation, hydrodynamic models, originally thought to be valid only for large systems such as $AA$ collisions, have proven quantitatively successful in describing several low-momentum observables in small systems as well [8, 9, 10, 11] (a non-exhaustive list of alternative explanations includes [12, 13, 14, 15, 16, 17]).
This hydrodynamic picture requires the presence of a deconfined quark-gluon plasma (QGP), implying final-state interactions and consequently jet quenching [18, 19, 20]. In $AA$ collisions, jet modifications are firmly established, and their dependence on the medium size is consistent with expectations from parton energy-loss calculations performed using perturbative QCD or holographic techniques [21, 22, 23, 24]. However, in $pA$ collisions, jet quenching remains elusive, possibly due to the short path lengths involved [25, 26, 27, 28]. Despite the absence of measurable energy loss, significant azimuthal anisotropies have been observed in high-$p_{T}$ particles in $pA$ systems [29, 30, 31]. In $AA$ collisions, this anisotropy is understood in terms of differences in energy loss due to the different path lengths traversed as a function of the jet orientation with respect to the reaction plane of the collision [32, 33], adding further puzzling elements to the picture we attempt to draw for small systems.
Recent LHC runs on OO and NeNe collisions have provided a valuable bridge between small ($pp$ and $pA$) and large (such as PbPb and AuAu) collision systems, thereby reducing uncertainties related to short path lengths and centrality determination [34, 35, 36, 37]. In these smaller systems, fluctuations and out-of-equilibrium dynamics are believed to play an increasingly critical role in achieving quantitative precision [38, 39, 40, 41].
In this work, we extend our semi-analytic jet quenching framework [42, 43, 44] by incorporating these elements to enable reliable predictions for the upcoming experimental results on the potential existence of jet quenching effects in OO collisions. We account for medium fluctuations by coupling the framework to state-of-the-art event-by-event hydrodynamic simulations of the expanding fireball within which jets are produced and subsequently traverse. Before hydrodynamics is believed to be applicable, we incorporate quenching effects during the pre-equilibrium phase via the hydrodynamic attractor extracted from QCD kinetic theory, representing the first time this is done in a realistic model for jet quenching.
We benchmark our model against a broad set of $AA$ data, including both LHC and RHIC energies and collision systems, spanning wide ranges in centrality, jet cone size and transverse momentum. Using Bayesian parameter estimation, we achieve a simultaneous and consistent description of jet suppression $R^{\rm jet}_{AA}(p_{T})$ and elliptic anisotropy $v_{2}^{\rm jet}(p_{T})$ across all available data. We show that the combination of $R^{\rm jet}_{AA}$ and $v^{\rm jet}_{2}$ provides constraints on energy-loss effects in the pre-equilibrium phase. The resulting posterior distributions can then be used to predict jet and hadron observables in other collision systems, including OO as we do in the present work.
II Framework
Our jet quenching framework builds on previous results from [42, 43, 44], which we briefly summarize here. We introduce two key new components: (i) the coupling to event-by-event fluctuating hydrodynamics, and (ii) energy loss in the pre-equilibrium phase using the hydrodynamic attractor.
Our starting point is the vacuum jet cross section, $\sigma^{pp}_{i}(p_{T},R)$, computed using MadGraph+Pythia [45, 46, 47] at NLO+LL accuracy. We reconstructed jets and assigned their flavor ($i=q/g$) with the Flavor-kt [48]. Nuclear PDF (nPDF) effects are incorporated using EPPS21 [49] within the LHAPDF [50] framework, yielding the modified cross section $\tilde{\sigma}^{pp}(p_{T},R)$. The charged hadron spectrum is evaluated by convoluting this jet spectrum with the NPC23 [51] fragmentation function (FF).
The quenched jet cross section is described via the convolution of the nPDF-modified jet cross section with the energy-loss probability distribution:
$\begin{split}\sigma^{AA}(p_{T},R)&=\sum_{i=q,g}\int d\varepsilon P_{i}(\varepsilon,R,p_{T})\tilde{\sigma}^{pp}_{i}(p_{T}+\varepsilon,R)\,.\end{split}$ (1)
Energy is lost predominantly due to medium-induced emissions escaping the jet cone $R$, as encoded in $P_{i}(\varepsilon,R,p_{T})$. The emission rate is computed using the Improved Opacity Expansion in a static medium, which captures both single hard and multiple soft scatterings, and also includes finite path length corrections [52, 53]. Medium properties affect these rates through LO Hard Thermal Loop results, namely the (bare) jet quenching parameter $\hat{q}_{0}=g^{2}_{\rm med}N_{c}m_{D}^{2}T/(4\pi)$, and the Debye screening mass $m_{D}^{2}=3g_{\rm med}^{2}T^{2}/2$, where $T$ is the local temperature of the hydrodynamic medium. Here, $g_{\rm med}$ is the (fixed, i.e. non-running) strong coupling for jet-medium interactions, representing the first of the two parameters we will estimate using Bayesian inference. We assume factorization between induced collinear emissions and transverse momentum broadening: semi-hard emissions undergo Gaussian broadening, while hard emissions inherit their transverse momentum from the Coulomb tail of the scattering potential. Since the softest emissions thermalize rapidly [54], they should be considered part of the medium and contribute to hydrodynamic wakes [55, 56], a manifestation of medium response to the jet passage. We approximate these physics by assuming that the thermalized energy is distributed uniformly around the jet hemisphere. Elastic energy loss is also included and constrained via the Einstein relation $\hat{e}=\hat{q}/(4T)$, and is assumed to thermalize in a similar fashion. Here, $\hat{q}$ accounts for the logarithmic corrections of $\hat{q}_{0}$ (see later).
Energetic jets consist of multiple partons produced by the high-virtuality DGLAP evolution that occurs at the early stages of the in-medium jet propagation [57]. These partons represent an additional source of energy loss. We account for this by resumming these multiple sources via a factorization between early vacuum-like and medium-induced emissions [58]. Importantly, we also include color-coherence, and -resolution [59, 60, 61] effects (referred to as coherent energy loss) by incorporating the non-local angular scale below which the medium cannot resolve individual partons, $\theta_{c}\sim 1/\sqrt{\hat{q}L^{3}}$, with $L$ being the traversed medium length. Detailed formulas are provided in [44].
Jet propagation occurs within an event-by-event fluctuating 2+1D hydrodynamic background. We employ hydro profiles from the comprehensive multi-stage framework IP-Glasma+JIMWLK+MUSIC+UrQMD, which successfully describes a wide range of bulk observables in heavy-ion collisions [62]. This state-of-the-art framework evolves the bulk starting from the initial energy density deposited at specific hotspots determined by the IPSat [63]. Jet probes are placed precisely onto these hotspots, whose number is proportional to the number of binary collisions $N_{\rm coll}$, and their initial orientation is uniformly distributed in the azimuthal angle $\phi$. Straight jet trajectories are sampled from their production point until they exit the deconfined medium at $T<T_{c}$, with $T_{c}$ marks the crossover transition to the hadron gas phase. The medium and jet-medium interaction quantities needed for the computation of radiative and elastic energy loss are then obtained by averaging along each jet trajectory, as extracted from the hydrodynamic profiles. We neglect further quenching below $T_{c}$. For more details on the embedding of our jet quenching framework in the realistic heavy-ion environment, see Ref. [44], where event-averaged hydro profiles were used instead. Event classification into centrality classes is based on the final charged particle multiplicity of the bulk, from which one also extracts the soft event-plane angles $\Psi_{n}$, necessary for computing the jet flow coefficients. Thus, our framework predicts the quenched jet spectrum for various parameters $\sigma^{AA\%}(p_{T},\eta,\phi,R)$.
The hydrodynamic description of the bulk begins at a finite time $\tau_{\rm hyd}=0.4$ fm. The initial condition of this nearly ideal fluid is reached through the Glasma evolution of saturated gluon fields (for reviews on this approach to thermalization, see [64, 65]). Quenching effects have traditionally been neglected before this time, largely due to the absence of reliable energy-loss calculations in the non-equilibrated phase. While an earlier start could be mimicked by adjusting the jet-medium coupling, combined observables such as high-$p_{T}$ $R_{AA}$ and $v_{2}$ are known to be sensitive to the actual value of the quenching onset time [66, 67]. The need to address jet quenching in the earliest stages has been emphasized in recent works where the pre-hydro phase is modeled either by Glasma [68, 69] or by QCD effective kinetic theory (EKT) [70, 71]. Despite differences in approach, these studies consistently conclude that the broadening per unit length in the short pre-equilibrium phase is considerably larger than in the equilibrated hydro phase.
In the present work, we extend our energy loss framework to the pre-equilibrium stage as follows. The out-of-equilibrium evolution of the local energy density has been extensively studied in EKT [72, 73, 74, 75, 76, 77], and it is well captured by the hydrodynamic attractor. We use the attractor to extrapolate the effective local temperature for times $\tau$ earlier than $\tau_{\rm hyd}$ as [78, 79]:
[…]
$(\tau^{1/3}T)_{\rm hyd}=\tau^{1/3}_{\rm hyd}\left(T(x)+\frac{2}{3}\frac{\eta/s}{\tau_{\rm hyd}}\right)\,$, (3)
with $\eta/s=0.12$, where $T(x)$ is the local initial temperature. We assume that our energy-loss model can be applied without any additional modification by using these extrapolated pre-equilibrium temperatures. Before $\tau_{\rm min}$, no quenching is applied, and we take this time as the second parameter of our Bayesian inference.
Finally, fluid velocity fields also evolve during the pre-equilibrium phase, as initial spatial anisotropies begin to generate flow [65]. This non-equilibrium evolution is well approximated by the universal pre-flow extracted from EKT [80]. We incorporate it by scaling the transverse velocity linearly with time:
[…]
where the longitudinal velocity scales with the space-time rapidity $\eta_{s}$.
III Results
We evaluate the quenched inclusive jet spectrum $\sigma^{AA\%}(p_{T},\eta,\phi,R)$ for various collision systems and centralities. Our focus is on the simultaneous description of $R^{\rm jet}_{AA}$ and $v^{\rm jet}_{n}$ observables, which we evaluate using
$\begin{split}R^{\rm jet}_{AA}=&\sigma^{AA\%}(p_{T},R)/\sigma^{pp}(p_{T},R)\,,\\ v^{\rm jet}_{n}=&\frac{\int d\eta d\phi\cos(n(\phi-\Psi_{n}))\sigma^{AA\%}(p_{T},\eta,\phi,R)}{\sigma^{AA}(p_{T},R)}\,.\end{split}$ (5)
where $\sigma(p_{T},R)=\int d\eta d\phi\sigma(p_{T},\eta,\phi,R)$ with the appropriate experimental cuts. We use the event plane method to compute $v_{n}^{\rm jet}$, following the procedure in the ATLAS measurement to which we compare our model [81]. We evaluate the hadron spectrum as
$\sigma(p^{h}_{T},\eta,\phi)=\sum_{i=q,g}\int^{1}_{x}\frac{{\rm d}z}{z}\sigma_{i}(\tfrac{p^{h}_{T}}{z},R,\eta,\phi)\,D_{i}^{h}(z,\mu_{F}^{2})\,$, (6)
where $x=2p_{T}^{h}\cosh(\eta)/\sqrt{s}$, and we used the jet scale $\mu_{F}=p_{T}^{h}R/z$. We averaged over $(u,d,s)$ FFs and neglected the others. This simple convolution neglects medium effects in fragmentation, and thus ignores, for example, the impact of the angular redistribution of particles due to broadening.
Figure 1: Parameter posterior probability distributions and correlations. Red distributions include $R^{\rm jet}_{AA}$ and $v_{2}^{\rm jet}$ data, while blue only includes $R_{AA}^{\rm jet}$.
Our work resembles previous studies of $R_{AA}$ and $v_{2}$ [82, 83, 66, 84, 85, 86, 87, 88, 89, 67, 90, 91, 92, 93, 94, 95, 96]. In contrast to these works, we focus on jets instead of hadrons, since jets are infrared- and collinear-safe, and have small non-perturbative corrections in vacuum [97, 98, 99, 100]. We use state-of-the-art hydrodynamics and energy-loss formulas, including elastic energy loss as well as the effects of medium response. To our knowledge, this is the first model implementation of jet energy loss in the pre-equilibrium phase carried out in a physically well-motivated manner.
Figure 2: Nuclear modification for jets and charged hadrons in OO collisions for two different centrality selections. Dark bands represent the 68% confidence interval for quenching parameters, and light bands represent the 68% nPDF uncertainties.
We employ Bayesian inference, implemented via the JETSCAPE/STAT toolkit [101, 102, 103], to constrain the jet-medium strong coupling $g_{\rm med}$ and the initial quenching time in the pre-equilibrium phase $\tau_{\rm min}$. We used flat priors in the intervals $1.2\leq g_{\rm med}\leq 1.7$, and $0.01\leq\tau_{\rm min}\,[{\rm fm}]\leq 0.4$. Our data points include $R^{\rm jet}_{AA}$ at LHC energies for different jet-cones and centrality selections [104, 105, 106], as well as jet $v_{2}^{\rm jet}$ [81], up to 40-50% centrality. The inference incorporates statistical and systematic uncertainties of the data, and we estimate the full correlation matrix by employing a finite correlation length of 0.2 [102, 107]. Luminosity and centrality uncertainties are added to the systematics. We neglect inherent model uncertainties in the inference such as scale and nPDF uncertainties, and we will present them separately for the OO predictions.
Figure. 1 shows the parameter posterior for two different scenarios. Fitting only $R^{\rm jet}_{AA}$ (blue) results in little constraint on $\tau_{\rm min}$. This is expected, as an earlier starting time can be compensated by a relatively smaller medium coupling to produce the same energy loss. Including $v_{2}^{\rm jet}$ constrains $\tau_{\rm min}\approx 0.24$ fm (red). This inference illustrates the role of pre-equilibrium effects in $v^{\rm jet}_{2}$.
Before moving to our predictions for oxygen collisions, Fig. 4 in App. A shows the posterior predictive distributions. We observe excellent overall agreement with all data. Small tensions can be found in $v_{2}^{\rm jet}$ at low $p_{T}$, which motivates further improvements of the energy loss framework. For completeness, we also show inference results on $g_{\rm med}$ with two limiting scenarios: early pre-equilibrium with fixed $\tau_{\min}=0.08$ fm, and no pre-equilibrium with $\tau_{\min}=0.4$ fm. The best fit (whose parameters are shown in Fig. 1) outperforms these two scenarios, although differences are not very significant. The extracted bare jet transport coefficient is approximately $\hat{q}_{0}/T^{3}\approx 1.7$, while after logarithmic corrections the effective value is $\hat{q}/T^{3}=\hat{q}_{0}/T^{3}\log(Q_{s}^{2}/\mu_{*}^{2})\approx 7$, where $Q_{s}^{2}=\hat{q}_{0}L$, and $\mu_{*}^{2}=m_{D}^{2}e^{-2+2\gamma_{E}}/4$. We note that this extracted $\hat{q}$ is consistent with other recent extractions using Bayesian inference in the JETSCAPE framework [103].
Figure 5 in App. A provides predictions for jets at RHIC energies [108] and charged hadrons at LHC energies [109, 110, 111, 112] using the extracted parameters. Even though our framework was designed for jets, it also provides a good description of the charged hadron $R_{AA}$ without including them in the inference. We also see, however, that our framework underestimates hadron elliptic flow at low $p_{T}$. This failure is consistent with previous findings using pQCD energy loss [66, 67]. It has been pointed out that a rather late quenching time of around $\sim 1$ fm would resolve the disagreement [67]. Alternatively, the absence of this issue for jets signals the need for model improvements which could be more relevant for (mid-$p_{T}$) hadrons than for jets. Again, for completeness, we show predictions with the previous two limiting scenarios (early pre-eq. and no pre-eq. energy loss), resulting as before in subtle differences.
OO predictions:
The recent OO run at the LHC has attracted considerable attention, with numerous dedicated studies on, for example, baseline analyses [113, 114, 115, 116] and hadron quenching predictions [117, 118, 119, 120, 121, 122, 123, 124, 125]. Compared to these studies, our framework incorporates additional improvements, with an emphasis on jet observables. For our OO predictions of $R_{AA}$ and $v_{2}$ we have set $\sqrt{s_{NN}}=5.36$ TeV and $|\eta|<2.1$. We defer to future work the computation of higher-order flow harmonics, along with the corresponding predictions for NeNe collisions.
Figure 2 shows our jet and charged hadron predictions for OO collisions. The left panel shows the inclusive jet suppression, where the no-quenching baseline, $\tilde{\sigma}^{pp}/\sigma^{pp}$, is indicated with gray bands. The dark band represents NLO scale uncertainties, while the light band is the 68% confidence of the nPDF. Below 50 GeV for jets and 30 GeV for hadrons, we underestimate nPDF uncertainties because of the relatively high momentum cutoff we chose for the NLO cross-sections [115, 116]. This cutoff does not affect the central value. Quenching effects are shown with red and blue for different centrality selections. Dark bands are the 68% confidence of our Bayesian parameter estimation, and on top of these, 68% nPDF uncertainties are shown with light bands. Peripheral collisions (30-100%) are less suppressed than central ones (0-30%), as expected, and clear quenching signals are present towards lower jet $p_{T}$ for both centrality selections. The right panel in Fig. 2 shows the charged hadron suppression, which presents a similar pattern to jets.
Figure 3 shows the inclusive jet and charged hadron elliptic flow. There is relatively smaller sensitivity to the nPDF, as $v_{2}$ is an observable that is differential in the transverse plane, using solely the medium spectra, so the impact of the nPDF largely cancels in Eq. (5). Semi-inclusive observables can present analogous advantages [115]. We observe that, in contrast to large collision systems, the elliptic flow coefficient $v_{2}$ increases with centrality. This is a consequence of the more prominent role played by the fluctuations of the initial nuclear shape in small collision systems [117, 126, 119]. Jet distributions are not shown for very low $p_{T}$, where our framework is not expected to apply.
Due to the relatively small path lengths involved in OO, the typical value of the coherence angle is relatively large, $\theta_{c}\sim 0.5$, and so $R=0.4$ jets are unresolved and lose energy as single color charges. Therefore, hadron and jet suppression in our framework can be understood by the mere $p_{T}$ shift resulting from hadronization. In contrast, in larger systems (with smaller values of $\theta_{c}$) both $R_{AA}$ and $v_{2}$ receive additional contributions due to resolved jet substructure fluctuations [127, 44].
Finally, Fig. 6 in App. B shows our OO predictions for minimum bias with the different limiting cases introduced previously. Only subtle differences are observed for earlier vs. later quenching for the ranges of $\tau_{\rm min}$ explored.
Figure 3: Charged hadron and jet elliptic flow in oxygen collisions for two different centrality selections. Dark bands represent the 68% confidence interval for quenching parameters, while light bands represent the 68% nPDF uncertainties.
IV Conclusions
In this work, we have integrated our semi-analytical jet quenching model with a state-of-the-art description of heavy-ion collisions, IP-Glasma+JIMWLK+MUSIC+UrQMD. We incorporated energy loss in the pre-equilibrium phase via the hydrodynamic attractor for the first time. Our framework provides a simultaneous description of jet $R_{AA}$ and $v_{2}$ in agreement with data in $AA$ collisions, offering a strong validation of the approach. We show that a joint description of these observables places constraints on the onset time of jet-medium interactions, which is around 0.2 fm, deep in the pre-equilibrium phase. Then, we applied these parameters to predict jet and hadron quenching in other collision systems. We saw that our framework successfully predicts hadron suppression but underestimates hadron elliptic flow towards lower $p_{T}$. Previous studies have reported similar results, suggesting that a delayed onset of quenching of about 1 fm could account for the hadron experimental data. Our success in describing jets, however, would rather suggest the need to account for additional medium effects whose importance could grow at lower $p_{T}$.
Our constrained framework has been used to make realistic predictions for the intermediate-sized OO collisions. We obtained sizable energy loss both for jets and charged hadrons, in quantitative agreement with the preliminary results from CMS [128]. Even though we have shown that pre-equilibrium effects play a moderate role in our model, we anticipate that including forthcoming data on high-$p_{T}$ flow coefficients will result in further constraints. Furthermore, in this work we have provided predictions for elliptic anisotropies and different centrality selections, both for hadrons and jets. Due to the small path lengths in OO, unlike in PbPb or AuAu, jet constituents are barely resolved (jets are fully coherent), and thus we obtain similar energy loss for hadrons and jets (modulo a $p_{T}$ shift in fragmentation). This outcome constitutes a prediction of our framework which can be confronted with future experimental results on jet suppression in light ions.
As an outlook, our jet quenching framework could benefit from an improved description of medium-induced emissions, elastic scatterings, and medium response in several ways. It is important to note that, at very early times, local color fields are saturated and highly anisotropic, a scenario which our energy-loss framework, built on local equilibrium, cannot accommodate. For these initial stages, it would be desirable to incorporate recent developments in energy-loss calculations for anisotropic media [129, 130, 131, 132, 133, 134], correlated color fields [135], and out of-equilibrium plasma [136, 137], so as to extend quenching for those earliest times where the attractor solution is not meant to apply.
Acknowledgements.
We are grateful to Björn Schenke and Chun Shen for sharing the hydrodynamic profiles and to Jannis Gebhard for sharing oxygen events. We greatly appreciate discussions with Aleksas Mazeliauskas, Yacine Mehtar-Tani, and Konrad Tywoniuk. The work of A.T. is supported by DFG through Emmy Noether Programme (project number 496831614) and through CRC 1225 ISOQUANT (project number 27381115). D.P. is funded by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101155036 (AntScat), by the European Research Council project ERC-2018-ADG-835105 YoctoLHC, by the Spanish Research State Agency under project PID2020-119632GB-I00, by Xunta de Galicia (CIGUS Network of Research Centres) and the European Union, and by Unidad de Excelencia María de Maetzu under project CEX2023-001318-M. D.P. also acknowledges support from the the Ramón y Cajal fellowship RYC2023-044989-I.
Appendix A Bayesian parameter estimation
We have already introduced our Bayesian framework in Sec. II, and here we extend this discussion by providing additional details. In this appendix, PbPb and AuAu cross-sections $\tilde{\sigma}^{AA}(p_{T},\eta,R)$ are computed using the microjet evolution as employed in [42, 44], with initial conditions from LO Pythia at $R_{0}=1$. Our framework is built on the JETSCAPE/STAT toolkit [101, 102, 103]. We evaluated our model for 24 design points, equally distributed in the parameter space. To focus on primary model features that are robust against statistical uncertainty, we reduced our data to its 10 principal components, and trained a Gaussian Process Emulator to interpolate between the predictions. The resulting parameter posterior is shown in the main text in Fig. 1. Furthermore, we provide two alternative inferences for $g_{\rm med}$: one with fixed $\tau_{\min}=0.08$ fm, (and extracted $\langle g_{\rm med}\rangle=1.385$) referred to as “early preeq”, and another with $\tau_{\min}=0.4$ fm, (extracted $\langle g_{\rm med}\rangle=1.473$), referred to as “no preeq”.
Figure 4 shows the corresponding posterior predictive distributions. These jet data were used to constrain the model parameters. Different rows correspond to different measurements. We observe good agreement with all data for both $R^{\rm jet}_{AA}$ and $v^{\rm jet}_{2}$, including all collision energies, centralities, rapidities, jet cone sizes, and momenta. There is a slight tension in $v^{\rm jet}_{2}$ for the lowest $p_{T}$ bins. This tension could be do to missing ingredients in the description of energy loss towards lower $p_{T}$. Dotted and dashed lines show the two alternative fits, where $\tau_{\min}$ is kept fixed. There are subtle differences between different fits, with the fixed earlier quenching reducing the jet $v_{2}$, as expected.
Figure 4: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ and $v_{2}$ data across all LHC experiments.
Finally, we also compare our results to jets at RHIC energies and charged hadron observables, shown in Fig. 5. As our framework was designed for jets, we did not include these data in the inference. We see an overall good agreement with both RHIC energies and charged hadron $R_{AA}$. Interestingly, hadron $R_{AA}$ seems to favor early quenching, as indicated by the dashed lines. There is a clear underestimation of the hadron elliptic flow at low momenta, consistent with previous findings from other theory collaborations, likely linked to the (milder) tension with low $p_{T}$ jets mentioned above.
Figure 5: Posterior predictive distributions for RHIC energy and for hadrons (not included in the inference). Data shown for RHIC jets and charged hadrons across different LHC experiments.
Appendix B OO predictions without pre-equilibrium quenching
Figure 6 shows the nuclear modification and elliptic flow for minimum bias OO collisions. Curves represent the limiting scenarios tested in App. A regarding pre-equilibrium quenching: “early preeq” denotes $\tau_{\min}=0.08$ fm, “no preeq” $\tau_{\min}=0.4$ fm, with their corresponding inferred $g_{\rm med}$. The best fit denotes our default setting shown in Fig. 1. For simplicity, we don’t show nPDF uncertainties. Slight differences are present in the different limiting scenarios. Interestingly, an earlier quenching can result in a relatively larger $v_{2}$, in contrast to larger systems, likely reflecting the more important role of fluctuations over average geometry in smaller systems.
Figure 6: Nuclear modification and elliptic flow for jets and charged hadrons in minimum bias OO collisions. Different lines corresponds to different limiting scenarios in the pre-equilibrium quenching. The bands represent the 68% confidence interval for quenching parameters.

### v2

@@
 The quenched jet cross section is described via the convolution of the nPDF-modified jet cross section with the energy-loss probability distribution:
-$\begin{split}\sigma^{AA}(p_{T},R)&=\sum_{i=q,g}\int d\varepsilon P_{i}(\varepsilon,R,p_{T})\tilde{\sigma}^{pp}_{i}(p_{T}+\varepsilon,R)\,.\end{split}$ (1)
+$split\sigma^{AA}(p_{T},R)&=\sum_{i=q,g}\int d\varepsilon P_{i}(\varepsilon,R,p_{T})\tilde{\sigma}^{pp}_{i}(p_{T}+\varepsilon,R)\,.$ (1)
 Energy is lost predominantly due to medium-induced emissions escaping the jet cone $R$, as encoded in $P_{i}(\varepsilon,R,p_{T})$.
-The emission rate is computed using the Improved Opacity Expansion in a static medium, which captures both single hard and multiple soft scatterings, and also includes finite path length corrections [52, 53].
+The emission rate is computed using the Improved Opacity Expansion in a static medium, which captures both single hard and multiple soft scatterings, and also includes finite path length corrections [52, 53] [^1] The static medium rates can mimic the rates of expanding medium in the multiple soft scattering limit [54, 55]..
 Medium properties affect these rates through LO Hard Thermal Loop results, namely the (bare) jet quenching parameter $\hat{q}_{0}=g^{2}_{\rm med}N_{c}m_{D}^{2}T/(4\pi)$, and the Debye screening mass $m_{D}^{2}=3g_{\rm med}^{2}T^{2}/2$, where $T$ is the local temperature of the hydrodynamic medium.
@@
 We account for this by resumming these multiple sources via a factorization between early vacuum-like and medium-induced emissions [60].
-Importantly, we also include color-coherence, and -resolution [59, 60, 61] effects (referred to as coherent energy loss) by incorporating the non-local angular scale below which the medium cannot resolve individual partons, $\theta_{c}\sim 1/\sqrt{\hat{q}L^{3}}$, with $L$ being the traversed medium length.
+Importantly, we also include color-coherence, and -resolution [61, 62, 63] effects (referred to as coherent energy loss) by incorporating the non-local angular scale below which the medium cannot resolve individual partons, $\theta_{c}\approx 1/\sqrt{\hat{q}L^{3}}$, with $L$ being the traversed medium length.
 Detailed formulas are provided in [44].
@@
 The medium and jet-medium interaction quantities needed for the computation of radiative and elastic energy loss are then obtained by averaging along each jet trajectory, as extracted from the hydrodynamic profiles.
-We neglect further quenching below $T_{c}$.
+For example, we evaluate the Debye mass as $m_{D}^{2}\propto(\int{\rm d}x_{F}T^{2}(x))/L$ for a given trajectory, where $x_{F}$ is the distance traversed in the fluid rest frame, thereby accounting for different flow velocities, and $L$ is the integral of ${\rm d}x_{F}$.
+The expressions for other relevant quantities can be found in Eq.
+(29) in [44].
+We neglect further quenching below the freeze-out temperature $T_{c}$.
 For more details on the embedding of our jet quenching framework in the realistic heavy-ion environment, see Ref.
@@
 The hydrodynamic description of the bulk begins at a finite time $\tau_{\rm hyd}=0.4$ fm.
-The initial condition of this nearly ideal fluid is reached through the Glasma evolution of saturated gluon fields (for reviews on this approach to thermalization, see [64, 65]).
-Quenching effects have traditionally been neglected before this time, largely due to the absence of reliable energy-loss calculations in the non-equilibrated phase.
+This parameter was fixed when the IP-Glasma+JIMWLK+MUSIC+UrQMD framework was tuned to LHC and RHIC data.
+The initial condition of this nearly ideal fluid is reached through the Glasma evolution of saturated gluon fields.
+The Glasma evolution should be followed by a kinetic evolution that drives the system toward hydrodynamics; however, this is not the setup used in the current bulk evolution framework.
+For reviews of this approach to thermalization, see [66, 67].
+Quenching effects have traditionally been neglected before $\tau_{\rm hyd}$, largely due to the absence of reliable energy-loss calculations in the non-equilibrated phases.
 While an earlier start could be mimicked by adjusting the jet-medium coupling, combined observables such as high-$p_{T}$ $R_{AA}$ and $v_{2}$ are known to be sensitive to the actual value of the quenching onset time [68, 69].
@@
 Despite differences in approach, these studies consistently conclude that the broadening per unit length in the short pre-equilibrium phase is considerably larger than in the equilibrated hydro phase.
-In the present work, we extend our energy loss framework to the pre-equilibrium stage as follows.
+In the present work, we extend our energy loss framework to the kinetic pre-equilibrium stage as follows.
 The out-of-equilibrium evolution of the local energy density has been extensively studied in EKT [74, 75, 76, 77, 78, 79], and it is well captured by the hydrodynamic attractor.
@@
 We assume that our energy-loss model can be applied without any additional modification by using these extrapolated pre-equilibrium temperatures.
-Before $\tau_{\rm min}$, no quenching is applied, and we take this time as the second parameter of our Bayesian inference.
-Finally, fluid velocity fields also evolve during the pre-equilibrium phase, as initial spatial anisotropies begin to generate flow [65].
+Before $\tau_{\rm min}$ , and for the Glasma phase, no quenching is applied, and we take this time as the second parameter of our Bayesian inference.
+Finally, fluid velocity fields also evolve during the kinetic pre-equilibrium phase, as initial spatial anisotropies begin to generate flow [67].
 This non-equilibrium evolution is well approximated by the universal pre-flow extracted from EKT [82].
@@
 Our focus is on the simultaneous description of $R^{\rm jet}_{AA}$ and $v^{\rm jet}_{n}$ observables, which we evaluate using
-$\begin{split}R^{\rm jet}_{AA}=&\sigma^{AA\%}(p_{T},R)/\sigma^{pp}(p_{T},R)\,,\\ v^{\rm jet}_{n}=&\frac{\int d\eta d\phi\cos(n(\phi-\Psi_{n}))\sigma^{AA\%}(p_{T},\eta,\phi,R)}{\sigma^{AA}(p_{T},R)}\,.\end{split}$ (5)
+$split R^{\rm jet}_{AA}=&\sigma^{AA\%}(p_{T},R)/\sigma^{pp}(p_{T},R)\,,\\ v^{\rm jet}_{n}=&\frac{\int d\eta d\phi\cos(n(\phi-\Psi_{n}))\sigma^{AA\%}(p_{T},\eta,\phi,R)}{\sigma^{AA}(p_{T},R)}\,.$ (5)
 where $\sigma(p_{T},R)=\int d\eta d\phi\sigma(p_{T},\eta,\phi,R)$ with the appropriate experimental cuts.
 We use the event plane method to compute $v_{n}^{\rm jet}$, following the procedure in the ATLAS measurement to which we compare our model [83].
+[^2] We also compared our results using the event plane method against the more experimentally common scalar product method, in which one correlates the jet momenta with the soft sector (see e.g.
+[84]), and found very little change in $v_{2}$, below $4\%$, consistent with what was shown by ATLAS [85].
 We evaluate the hadron spectrum as
-$\sigma(p^{h}_{T},\eta,\phi)=\sum_{i=q,g}\int^{1}_{x}\frac{{\rm d}z}{z}\sigma_{i}(\tfrac{p^{h}_{T}}{z},R,\eta,\phi)\,D_{i}^{h}(z,\mu_{F}^{2})\,$, (6)
+[…]
 where $x=2p_{T}^{h}\cosh(\eta)/\sqrt{s}$, and we used the jet scale $\mu_{F}=p_{T}^{h}R/z$.
@@
 This simple convolution neglects medium effects in fragmentation, and thus ignores, for example, the impact of the angular redistribution of particles due to broadening.
-Figure 1: Parameter posterior probability distributions and correlations. Red distributions include $R^{\rm jet}_{AA}$ and $v_{2}^{\rm jet}$ data, while blue only includes $R_{AA}^{\rm jet}$.
 Our work resembles previous studies of $R_{AA}$ and $v_{2}$ [86, 84, 68, 87, 88, 89, 90, 91, 92, 69, 93, 94, 95, 96, 97, 98, 99].
@@
 To our knowledge, this is the first model implementation of jet energy loss in the pre-equilibrium phase carried out in a physically well-motivated manner.
+Figure 1: Parameter posterior probability distributions and correlations. Red distributions include $R^{\rm jet}_{AA}$ and $v_{2}^{\rm jet}$ data, while blue only includes $R_{AA}^{\rm jet}$.
 Figure 2: Nuclear modification for jets and charged hadrons in OO collisions for two different centrality selections.
 Dark bands represent the 68% confidence interval for quenching parameters, and light bands represent the 68% nPDF uncertainties.
+Figure 3: Charged hadron and jet elliptic flow in oxygen collisions for two different centrality selections.
+Dark bands represent the 68% confidence interval for quenching parameters, while light bands represent the 68% nPDF uncertainties.
 We employ Bayesian inference, implemented via the JETSCAPE/STAT toolkit [104, 105, 106], to constrain the jet-medium strong coupling $g_{\rm med}$ and the initial quenching time in the pre-equilibrium phase $\tau_{\rm min}$.
 We used flat priors in the intervals $1.2\leq g_{\rm med}\leq 1.7$, and $0.01\leq\tau_{\rm min}\,[{\rm fm}]\leq 0.4$.
+We do not consider the scenario when $\tau_{\min}>\tau_{\rm hyd}$ as we believe delaying quenching beyond hydrodynamics is unphysical.
 Our data points include $R^{\rm jet}_{AA}$ at LHC energies for different jet-cones and centrality selections [107, 108, 109], as well as jet $v_{2}^{\rm jet}$ [83], up to 40-50% centrality.
@@
 This inference illustrates the role of pre-equilibrium effects in $v^{\rm jet}_{2}$.
-Before moving to our predictions for oxygen collisions, Fig.
-4 in App.
-A shows the posterior predictive distributions.
+[^3] It would be interesting to include $\tau_{\rm hyd}$ in the inference, since it could affect our results.
+This would require the simultaneous tuning of the soft and hard sectors, a task that we defer to future work.
+Before moving to our predictions for oxygen collisions, Figs.
+4-5 in App.
+A show the posterior predictive distributions.
 We observe excellent overall agreement with all data.
 Small tensions can be found in $v_{2}^{\rm jet}$ at low $p_{T}$, which motivates further improvements of the energy loss framework.
-For completeness, we also show inference results on $g_{\rm med}$ with two limiting scenarios: early pre-equilibrium with fixed $\tau_{\min}=0.08$ fm, and no pre-equilibrium with $\tau_{\min}=0.4$ fm.
+For completeness, Figs.
+4-5 in App.
+A also show two additional inference results on $g_{\rm med}$ with two limiting scenarios: early pre-equilibrium with fixed $\tau_{\min}=0.08$ fm, and no pre-equilibrium with $\tau_{\min}=0.4$ fm.
 The best fit (whose parameters are shown in Fig.
@@
 We note that this extracted $\hat{q}$ is consistent with other recent extractions using Bayesian inference in the JETSCAPE framework [106].
-Figure 5 in App.
-A provides predictions for jets at RHIC energies [108] and charged hadrons at LHC energies [109, 110, 111, 112] using the extracted parameters.
+Figures 6-7 in App.
+A provide predictions for jets at RHIC energies [111] and charged hadrons at LHC energies [112, 113, 85, 114] using the extracted parameters.
 Even though our framework was designed for jets, it also provides a good description of the charged hadron $R_{AA}$ without including them in the inference.
@@
 This failure is consistent with previous findings using pQCD energy loss [68, 69].
-It has been pointed out that a rather late quenching time of around $\sim 1$ fm would resolve the disagreement [67].
+It has been pointed out that a rather late quenching time of around $\approx 1$ fm would resolve the disagreement [69].
 Alternatively, the absence of this issue for jets signals the need for model improvements which could be more relevant for (mid-$p_{T}$) hadrons than for jets.
@@
 2 shows the charged hadron suppression, which presents a similar pattern to jets.
-Figure 3 shows the inclusive jet and charged hadron elliptic flow.
-There is relatively smaller sensitivity to the nPDF, as $v_{2}$ is an observable that is differential in the transverse plane, using solely the medium spectra, so the impact of the nPDF largely cancels in Eq.
-(5).
-Semi-inclusive observables can present analogous advantages [115].
-We observe that, in contrast to large collision systems, the elliptic flow coefficient $v_{2}$ increases with centrality.
-This is a consequence of the more prominent role played by the fluctuations of the initial nuclear shape in small collision systems [117, 126, 119].
-Jet distributions are not shown for very low $p_{T}$, where our framework is not expected to apply.
-Due to the relatively small path lengths involved in OO, the typical value of the coherence angle is relatively large, $\theta_{c}\sim 0.5$, and so $R=0.4$ jets are unresolved and lose energy as single color charges.
+[…]
+Due to the relatively small path lengths involved in OO, the typical value of the coherence angle is relatively large, $\theta_{c}\approx 0.5$, and so $R=0.4$ jets are unresolved and lose energy as single color charges.
 Therefore, hadron and jet suppression in our framework can be understood by the mere $p_{T}$ shift resulting from hadronization.
@@
 later quenching for the ranges of $\tau_{\rm min}$ explored.
-Figure 3: Charged hadron and jet elliptic flow in oxygen collisions for two different centrality selections.
-Dark bands represent the 68% confidence interval for quenching parameters, while light bands represent the 68% nPDF uncertainties.
 IV Conclusions
@@
 We saw that our framework successfully predicts hadron suppression but underestimates hadron elliptic flow towards lower $p_{T}$.
-Previous studies have reported similar results, suggesting that a delayed onset of quenching of about 1 fm could account for the hadron experimental data.
-Our success in describing jets, however, would rather suggest the need to account for additional medium effects whose importance could grow at lower $p_{T}$.
+Our results provide a complementary perspective to previous phenomenological studies.
+By incorporating early-time energy loss, we achieve a simultaneous description of jet $R_{AA}$ and $v_{2}$, as well as hadron $R_{AA}$ and $v_{2}$ at high-$p_{T}$.
+This resolves the, in our opinion, unphysical delay of quenching invoked in earlier approaches (e.g.
+[68, 69]) to reproduce hadron $R_{AA}$ and $v_{2}$.
+The successful description of jet observables supports the onset of quenching at early times, whereas the remaining discrepancy in hadron $v_{2}$ at moderate transverse momentum indicates the necessity of additional medium effects beyond longitudinal energy loss.
 Our constrained framework has been used to make realistic predictions for the intermediate-sized OO collisions.
-We obtained sizable energy loss both for jets and charged hadrons, in quantitative agreement with the preliminary results from CMS [128].
+We obtained sizable energy loss both for jets and charged hadrons, in quantitative agreement with the results from CMS [130], where our predictions are also presented.
 Even though we have shown that pre-equilibrium effects play a moderate role in our model, we anticipate that including forthcoming data on high-$p_{T}$ flow coefficients will result in further constraints.
@@
 This outcome constitutes a prediction of our framework which can be confronted with future experimental results on jet suppression in light ions.
+All of our data and plots are available at [131].
 As an outlook, our jet quenching framework could benefit from an improved description of medium-induced emissions, elastic scatterings, and medium response in several ways.
 It is important to note that, at very early times, local color fields are saturated and highly anisotropic, a scenario which our energy-loss framework, built on local equilibrium, cannot accommodate.
-For these initial stages, it would be desirable to incorporate recent developments in energy-loss calculations for anisotropic media [129, 130, 131, 132, 133, 134], correlated color fields [135], and out of-equilibrium plasma [136, 137], so as to extend quenching for those earliest times where the attractor solution is not meant to apply.
+For these initial stages, it would be desirable to incorporate recent developments in energy-loss calculations for anisotropic and flowing media [132, 133, 134, 135, 136, 137, 138], correlated color fields [139], and out of-equilibrium plasma [140, 141], so as to extend quenching for those earliest times where the attractor solution is not meant to apply.
 Acknowledgements.
@@
 Furthermore, we provide two alternative inferences for $g_{\rm med}$: one with fixed $\tau_{\min}=0.08$ fm, (and extracted $\langle g_{\rm med}\rangle=1.385$) referred to as “early preeq”, and another with $\tau_{\min}=0.4$ fm, (extracted $\langle g_{\rm med}\rangle=1.473$), referred to as “no preeq”.
-Figure 4 shows the corresponding posterior predictive distributions.
+Figures 4-5 show the corresponding posterior predictive distributions.
 These jet data were used to constrain the model parameters.
@@
 There are subtle differences between different fits, with the fixed earlier quenching reducing the jet $v_{2}$, as expected.
-Figure 4: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ and $v_{2}$ data across all LHC experiments.
-Finally, we also compare our results to jets at RHIC energies and charged hadron observables, shown in Fig.
-5.
+Figure 4: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ and jet $v_{2}$ data from ATLAS.
+Figure 5: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ from ALICE and CMS.
+Finally, we also compare our results to jets at RHIC energies and charged hadron observables, shown in Figs.
+6-7.
 As our framework was designed for jets, we did not include these data in the inference.
@@
 There is a clear underestimation of the hadron elliptic flow at low momenta, consistent with previous findings from other theory collaborations, likely linked to the (milder) tension with low $p_{T}$ jets mentioned above.
+Figure 6: Posterior predictive distributions for RHIC energy and for hadrons (not included in the inference). Data shown for RHIC jets and charged hadrons across different LHC experiments.
 Figure 7: Posterior predictive distributions for RHIC energy and for hadrons (not included in the inference). Data shown for RHIC jets and charged hadrons across different LHC experiments.